OpenStudy (vespersaviour):

The value of y varies directly with x^2, and y = 64 when x = 4. What is the value of y when x = 6?

OpenStudy (vespersaviour):

y=kx^2 y=k(6 * 6) y=k(36) k=36 I believe now i took 36 * 6 to find y and it came out as 216. To test that i did 216 divided by 6 and it came out 36. Would y be 216?

OpenStudy (vespersaviour):

@YoungStudier can you help?

OpenStudy (youngstudier):

@rebeccaxhawaii

OpenStudy (vespersaviour):

I believe itd be 216 cause when doing the problem they gave y=64 x=4. 4*4 = 16 and 64 divided by 4 was 16... .3.

OpenStudy (photon336):

Take a look at this \[y = kx^{2}\] \[\frac{ y }{ x^{2} } = k \] \[\frac{ y }{ x^{2} } = \frac{ 64 }{ 4^{2} } = 4 = k \] \[y = 4*x^{2}\]

OpenStudy (photon336):

@VesperSaviour you need to find what k is first

OpenStudy (vespersaviour):

4^2 is 16 which should be K @Photon336

OpenStudy (photon336):

well, you know what y is and you know what x is right? so you need to solve for k \[\frac{ y }{x^{2} } = k \]

OpenStudy (vespersaviour):

if its 64 divided by 4^2 (16) then its 4...

OpenStudy (photon336):

Yes, make sure you re-arrange the formula for the unknown variable FIRST. then plug in the numbers.

OpenStudy (vespersaviour):

For the onewe gotta figure out... y=kx^2 x=6 so 6^2 = 36....from what you are saying we take the y divide it by the x for the k but we need to find y...i took 216 divided it by 36 and that came out as 6....i still think 216.. @Photon336

OpenStudy (photon336):

okay k =4 and x = 6 right?

OpenStudy (photon336):

\[4*(6)^{2} = y \]

OpenStudy (vespersaviour):

144?

OpenStudy (vespersaviour):

so for the y i have to figure out its 144? @Photon336

OpenStudy (photon336):

yep that's your y value

OpenStudy (vespersaviour):

thank you